# OptiCarVis: Improving Automated Vehicle Functionality Visualizations Using Bayesian Optimization to Enhance User Experience

### v1

@@ 3. OptiCarVis: Optimizing Automated Vehicle Functionality Visualizations
Visualizations on WSDs use icons (Lindemann et al., 2018), highlighting (Colley et al., 2021a), and other elements (e.g., see (Kunze et al., 2019)) to communicate AV functionalities, which could be crucial for user acceptance (Colley et al., 2021a). These visualizations aim to increase end-users’ perceived safety and trust in AVs, their understanding of AV actions in various driving situations (e.g., unexpected stopping), and reduce the cognitive load when processing this (potentially overwhelming) information.
@@ 3.2. Objective Functions
Normalization is required before submitting these to the optimizer because the subjective metrics values have ranges based on 20-, 5-, or 7-point Likert scales. We transformed these six metrics into the $[-1,1]$ range. After the transformation, the cognitive load objective is a function to be maximized (a higher value means less load).
@@ 4.5.2. Objective Measures
We recorded the Bayesian optimizer’s performance metrics and the time taken in the Unity application for questionnaire responses (C2-C6) and custom design (C2 and C6). For RQ3, we embedded the webcam-based eye-tracker UnitEye [^2] Will be open-sourced – own development. in the study application to track areas of interest (AOIs): pedestrian, vehicle, traffic sign, pedestrian intention icon, occluded car, CAD-covered area sphere, and vehicle status HUD. Participants calibrated the eye-tracker before the study so that we could monitor their focus and attention during the design and the AV ride on the 33-second and 3-minute driving routes. The eye-tracking data was not used as an objective function of the Bayesian optimizer.
@@ 5.1.1. Data Analysis
The error bars represent bootstrap confidence intervals (i.e., mean_cl_boot).
We refrain from reporting the Pareto front graphically due to (1) the high number of parameter value combinations—even with only five discrete levels for our nine continuous design parameters (four $s$ and five $\alpha$ values, see Table 1), there are approximately $5^{9}$ possible combinations—and (2) the resulting challenges in visualization due to the high number of dimensions ($p_{1}\dots p_{16}$).
@@ 5.3. Qualitative Results
After the final exposure to the visualizations, 22 participants provided open feedback on the expectation, satisfaction, and interactivity themes (see Section 4.5.1). We analyzed this feedback using a structured two-phase process involving three authors. In the first phase, each author independently categorized the feedback into positive, negative, or suggestive sentiments across the three themes, summarizing key statements and insights. In the second phase, the authors collaboratively reviewed and finalized which feedback to include, ensuring consistent sentiment assignment and resolving disagreements.
@@ 6.6. Limitations and Future Work
Our work has been instrumental in assessing the application of MOBO to improve the user experience of AV functionality visualizations. However, limitations exist. The choice of algorithm, including the acquisition function and other parameters, might have influenced the study outcomes. As with any optimization approach, the selected hyperparameters can significantly impact the results. Along with this, the participant selection might introduce biases, which could affect the generalizability of our findings. Besides, a limitation was using a webcam-based eye-tracking method, which has inherent inaccuracies (Wisiecka et al., 2022). Nonetheless, we argue that the inaccuracies are a necessary trade-off, primarily because the technique offered novel insights into users’ reactions to the HITL process and the resulting designs that are otherwise infeasible for an online study.
@@
Interface clutter is a recognized issue (Colley et al., 2020, 2022b, 2021a). An interface attempting to display all potential visualization parameters could become overwhelming, potentially distracting users or obscuring essential information. In our study, while we aimed for comprehensive visualization, the risk of cluttering the interface remained evident.
Cultural differences influence user perceptions of AVs (Edelmann et al., 2021; Lanzer et al., 2020), for instance, the acceptance of AV maneuvers (Edelmann et al., 2021), also likely affecting their perception of AV functionality visualizations. Our study was limited to one culture. However, the OptiCarVis approach using HITL optimization could accommodate diverse user needs stemming from cultural context. To assess its effectiveness, future research should evaluate OptiCarVis with users from different cultures.
@@ 7. Conclusion
We will make the Bayesian optimizer, the Unity application, and the collected (anonymized) data available upon acceptance. This will include installation instructions and information on required 3rd party Unity assets. The Unity project supports novel application scenarios by including easily adaptable settings regarding MOBO, server-client infrastructure for the online study, webcam-based eye tracking, and AV driving behavior.
@@ Acknowledgements.
We thank all study participants. Additionally, we thank the 6th Summer School on Computational Interaction [^3] https://cixschool2022.cs.uni-saarland.de/ for providing significant insights into computational optimization. This work was supported by a German Academic Exchange Service (DAAD) fellowship.

### v3

@@ 3. OptiCarVis: Optimizing Automated Vehicle Functionality Visualizations
Visualizations on WSDs use icons (Lindemann et al., 2018), highlighting (Colley et al., 2021a; Jansen et al., 2024), and other elements (e.g., see (Kunze et al., 2019)) to communicate AV functionalities, which could be crucial for user acceptance (Colley et al., 2021a; Jansen et al., 2024). These visualizations aim to increase end-users’ perceived safety and trust in AVs, their understanding of AV actions in various driving situations (e.g., unexpected stopping) (Jansen et al., 2024), and reduce the cognitive load when processing this (potentially overwhelming) information.
@@ 3.2. Objective Functions
Normalization is required before submitting these to the optimizer because the subjective metrics values have ranges based on 20-, 5-, or 7-point Likert scales. We transformed these six metrics into the $[-1,1]$ range. After this, the cognitive load objective is a function to be maximized (a higher value means less load).
@@ 4.5.2. Objective Measures
We recorded the Bayesian optimizer’s performance metrics and the time taken in the Unity application for questionnaire responses (C2-C6) and custom design (C2 and C6). For RQ3, we embedded the webcam-based eye-tracker UnitEye (Wagner et al., 2024) in the study application to track areas of interest (AOIs): pedestrian, vehicle, traffic sign, pedestrian intention icon, occluded car, CAD-covered area sphere, and vehicle status HUD. Participants calibrated the eye-tracker before the study so that we could monitor their focus and attention during the design and the AV ride on the 33-second and 3-minute driving routes. The eye-tracking data was not used as an objective function of the Bayesian optimizer.
@@ 5.1.1. Data Analysis
The error bars represent bootstrap confidence intervals (i.e.,
mean_cl_boot). We refrain from reporting the Pareto front graphically due to (1) the high number of parameter value combinations—even with only five discrete levels for our nine continuous design parameters (four $s$ and five $\alpha$ values, see Table 1), there are approximately $5^{9}$ possible combinations—and (2) the resulting challenges in visualization due to the high number of dimensions ($p_{1}\dots p_{16}$).
@@ 5.3. Qualitative Results
After the final exposure to the visualizations, 22 participants provided open feedback on the expectation, satisfaction, and interactivity themes (see Section 4.5.1). We analyzed this feedback using a structured two-phase process involving three authors. In the first phase, each author independently categorized the feedback into positive, negative, or suggestive sentiments across the three themes, summarizing key statements and insights. The inter-rater agreement was assessed using Fleiss’ kappa, which yielded a value of 0.77, indicating substantial agreement (Landis and Koch, 1977). In the second phase, the authors collaboratively reviewed and finalized which feedback to include, ensuring consistent sentiment assignment and resolving disagreements.
@@ 6.6. Limitations and Future Work
Our work has been instrumental in assessing the application of MOBO to improve the user experience of AV functionality visualizations. However, limitations exist. The choice of algorithm, including the acquisition function and other parameters, might have influenced the study outcomes. As with any optimization approach, the selected hyperparameters can significantly impact the results. Along with this, the participant selection might introduce biases, which could affect the generalizability of our findings. Besides, a limitation was using a webcam-based eye-tracking method, which has inherent inaccuracies (Wisiecka et al., 2022). Nonetheless, we argue that the inaccuracies are a necessary trade-off because the technique offered novel insights into users’ reactions to the HITL process and the resulting designs that are otherwise infeasible for an online study.
@@
Interface clutter is a recognized issue (Colley et al., 2020, 2022b, 2021a). An interface attempting to display all potential visualization parameters could become overwhelming, potentially distracting users or obscuring essential information. In our study, while we aimed for comprehensive visualization, the risk of cluttering the interface remained.
Cultural differences influence user perceptions of AVs (Edelmann et al., 2021; Lanzer et al., 2020) (e.g., acceptance of AV maneuvers (Edelmann et al., 2021)), also likely affecting their perception of AV functionality visualizations. Our study was limited to one culture. However, the OptiCarVis approach using HITL optimization could accommodate diverse user needs stemming from cultural context. To assess its effectiveness, future research should evaluate OptiCarVis with users from different cultures.
@@ 7. Conclusion
We make the Bayesian optimizer (see https://github.com/Pascal-Jansen/Bayesian-Optimization-for-Unity), the Unity application upon request, and the collected (anonymized) data (see https://github.com/M-Colley/opticarvis-data) available. The Unity project supports novel application scenarios by including easily adaptable settings regarding MOBO, server-client infrastructure for the online study, webcam-based eye tracking (Wagner et al., 2024), and AV driving behavior.
@@ Acknowledgements.
We thank all study participants. Additionally, we thank the 6th Summer School on Computational Interaction [^2] https://cixschool2022.cs.uni-saarland.de/ for providing significant insights into computational optimization. This research was supported by a German Academic Exchange Service (DAAD) fellowship and conducted in the context of the DFG project ”Computational Optimization of In-Vehicle User Interface Design for Maximized Usability, Safety, and Trust.”